OpenStudy (anonymous):

for the given equation list the intercepts and test for symmetry... y=x^2-x-20 is it symmetric to the x-asix,y-axis, and origin???

OpenStudy (anonymous):

What do you think? Graph it first.

OpenStudy (anonymous):

The intercepts are (0,20) and (-4,0) and (5,0) The first one we obtain by putting x=0 and finding y=20. The x-intercepts we get them by solving \[ x^2 - x - 20= (x+4)(x-5)=0 \]

OpenStudy (anonymous):

The graph is not symmetric with respect to the x-axis, y-axis or the origin